2 ft Q2) Find the coordinates for the centroid for the plate shown. 1.ft Tat-21- -3 ft
Answer #1

Giver Plate : ny E 2 ft. Н 3 eft: →a A B eft Kft from the figure; C is the ceigin (al. y) = (0,0) let ze = Distance of cent3 3 →y, Hz = Distance of centroid of CoE from yoqaeis >> 213 = I ft. (towards Right) Y, = Distance of contraid of AHGB from Hfrom ५ equc Distance of centroid from ze-ares = y € y, A, t. Y₂ Azt YzAz Ait A2 TA3 * j = (0.5 x 1) + (1-5%6) + (1x4-5) Sput

Centroid of the plate is at a distance of x = 0.3478 ft. towards left of y-axis & at a distance of y = 1.2174 ft. above the x-axis as shown in figure below (point O) :-

y E P O= centroid (म, ५) = (0.3478, H 1. 2074 उ 2 PP 0 →म B. C D 211 Kaft उt. * *

​​​​

And distance of centroid from left most point (point A in figure )

= ==3-03478 = 2.6522 ==3-03478 = 2.65223-0.3478

= 2.6522 ft. from point A.

And above x-axis = 1.2174 ft.

Here is full step by step solution of the given problem.

I hope this will help you & you will like it.

Don't forget to give a thumbs up.

THANK YOU...!